# Bochner’s technique in Einstein’s non-symmetric geometry

### v1

@@ 3 Main results
Here, motivated by the structure of an almost Lie algebroid, we develop Bochner’s technique for NGT spaces.
@@ 3.1 The modified covariant derivative and bracket
and the difference tensor $K$, see (10), but generally $\nabla^{f}$ and $\nabla^{0}$ are not linear connections on $M$. Set $\nabla^{f}_{X}\,\psi=({f}X)\psi=d\psi({f}X)$ for the f{f}-gradient of $\psi\in C^{1}(M)$. The f{f}-derivative of an $(p,k)$-tensor $S$, where $p=0,1$, is the following $(p,k+1)$-tensor $\nabla^{f}S$ :
$(\nabla^{f}S)(Y,X_{1},\ldots,X_{k}):=\nabla^{f}_{Y}(S(X_{1},\ldots,X_{k}))-\sum\nolimits_{\,i=1}^{k}S(X_{1},\ldots,\nabla^{f}_{Y}X_{i},\ldots,X_{k}).$ (23)
Such $\nabla^{f}$ (as well as $\nabla^{0}$) satisfies Koszul conditions (with any $\psi\in C^{\infty}(M)$), see (21),
@@
We use the standard notation $\nabla^{f}_{Y}\,S=(\nabla^{f}S)(Y,\ldots)$. A tensor $S$ is called $\nabla^{f}$-parallel if $\nabla^{f}S=0$.
Using (23) and $\nabla^{g}\,g=0$, we obtain the following:
@@
For an f{f}-connection $\nabla^{f}$ on $(M,G=g+F)$, its dual f{f}-connection $\breve{\nabla}^{f}$ is defined by the equality
@@
If $k=1$, the last term in the above does not appear. If $k>1$, we fix $2\leq p\leq k$, and calculate
@@ 3.2 The f{f}-divergence and Bochner f{f}-Laplacian
using a local orthonormal frame $\{e_{i}\}$. The divergence of a vector field $X$ can also be defined by
@@
The “musical” isomorphisms $\sharp:T^{*}M\to TM$ and $\flat:TM\to T^{*}M$ are used for rank one tensors, e.g. if $\omega\in\Lambda^{1}(M)$ and $X\in{\mathfrak{X}}_{M}$ then $\omega(X)=g(\omega^{\sharp},X)=g(\omega,X^{\flat})=X^{\flat}(\omega^{\sharp})$.
In order to generalize the Stokes Theorem for distributions, we formulate the following.
@@
where $\nu$ is the unit inner normal to $\partial M$. In particular, on a manifold $(M,G=g+F)$ without boundary, for any $X\in\mathfrak{X}_{M}$ with compact support, (38) reduces to the equality $\int_{M}(\operatorname{div}_{f}X)d\operatorname{vol}_{g}{=}0$.
@@
while, for $k$-forms, we set
$g(\omega_{1},\,\omega_{2})=\sum\nolimits_{\,i_{1}<\ldots\,<i_{k}}\omega_{1}(e_{i_{1}},\ldots,\,e_{i_{k}})\,\omega_{2}(e_{i_{1}},\ldots,\,e_{i_{k}}).$
@@
The $L^{2}$-adjoint operator $\nabla^{*{f}}$ for $\nabla^{f}$ maps $(p,k+1)$-tensors, where $p=0,1$, to $(p,k)$-tensors:
$(\nabla^{*{f}}S)(X_{1},\ldots,X_{k})=-\sum\nolimits_{\,i}(\nabla^{f}_{e_{i}}S)(e_{i},X_{1},\ldots,X_{k}),$ (39)
and similarly for $\breve{\nabla}^{*{f}}$.
@@
Define a compactly supported 1-form $\omega$ by
$\omega(Y)=g(\iota_{\,Y}\,\omega_{2},\,\omega_{1}),\quad Y\in{\mathcal{X}}_{M}.$
To simplify calculations, assume that $k=1$. Take an orthonormal frame $(e_{i})$ such that $\nabla_{Y}e_{i}=0$ for $Y\in T_{x}M$. To simplify calculations, assume that $s=t=1$, then at $x\in M$,
$-\nabla^{*{f}}\omega$ $=\sum\nolimits_{j}(\nabla^{f}_{e_{j}}\omega)(e_{j})=\sum\nolimits_{i,j}g({f}e_{j},e_{i})\,(\nabla_{e_{i}}\omega)(e_{j})=\sum\nolimits_{i,j,c}g({f}e_{j},e_{i})\,\nabla_{e_{i}}g(\omega_{2}(e_{j},e_{c}),\omega_{1}(e_{c})\,)$
@@
The differential operator $\nabla^{*{f}}\,\nabla^{f}$ will be called the Bochner f{f}-Laplacian for an f{f}-connection.
@@
Let an Einstein’s connection $\nabla$ on a closed manifold $(M,G=g+F)$ satisfy (15) and (36). If $g(\nabla^{*f}\,\nabla^{f}\omega,\,\omega)\leq 0$ holds for a form $\omega\in\Lambda^{k}(M)$, then $\nabla^{f}\omega=0$, i.e., $\omega$ is f{f}-parallel.
@@
hence, $\nabla^{f}\omega=0$. ∎
@@ 3.3 The Hodge f{f}-Laplacians
Using an Einstein’s connection $\nabla$, we define the exterior f{f}-derivative of a form $\omega\in\Lambda^{k}(M)$ by
@@
The f{f}-derivative $d^{f}:{\mathrm{\Omega}}^{k}(M)\rightarrow{\mathrm{\Omega}}^{k+1}(M)$ is a 1-degree derivation, see (2.2), i.e.,
@@
+ $+\sum\nolimits_{\,0\leq i<j\leq k}(-1)^{i+j}\omega\big([X_{i},X_{j}]_{f},X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{j-1},X_{j+1},\ldots,X_{k}\big).$ (45)
@@
+ $+\sum\nolimits_{\,0\leq i<j\leq k}(-1)^{i+j}\omega\big([X_{i},X_{j}]^{\smile}_{f},X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{j-1},X_{j+1},\ldots,X_{k}\big).$ (46)
@@
+ $+\sum\nolimits_{\,0\leq i<j\leq k}(-1)^{i+j}\omega\big(\nabla^{f}_{X_{i}}X_{j}-\nabla^{f}_{X_{j}}X_{i},X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{j-1},X_{j+1},\ldots,X_{k}\big).$
@@
Define the f{f}-codifferential $\delta^{f}:\Lambda^{k}(M)\to\Lambda^{k-1}(M)$ by
$\delta^{\,f}:=\nabla^{*{f}}.$
Similarly, we set $\breve{\delta}^{\,f}:=\breve{\nabla}^{*{f}}$. Using the analog of (39), we obtain
@@
On a closed manifold $(M,G=g+F)$ with an Einstein’s connection $\nabla$ satisfying conditions (15) and (36), for any forms $\omega_{1}\in\Lambda^{k}(M)$ and $\omega_{2}\in\Lambda^{k+1}(M)$, we have
$(\breve{\delta}^{\,{f}}\omega_{2},\,\omega_{1})_{L^{2}}=(\omega_{2},\,d^{f}\omega_{1})_{L^{2}}.$
@@
$g(d^{f}\omega_{1},\,\omega_{2})$ $=\sum\nolimits_{\,u=0}^{k}(-1)^{i}\,\nabla^{f}_{\partial_{i_{u}}}\omega_{1}(\partial_{i_{0}},\ldots,\partial_{i_{u-1}},\partial_{i_{u+1}},\ldots,\partial_{i_{k}})g^{i_{0}j_{0}}\ldots g^{i_{k}j_{k}}\omega_{2}(\partial_{j_{0}},\ldots,\partial_{j_{k}})$
@@
A differential form $\omega$ is called f{f}-harmonic if ${\mathrm{\Delta}}_{H}^{\,{f}}\,\omega=0$ and $\|\omega\|_{L^{2}}<\infty$ (similarly for $\breve{\mathrm{\Delta}}^{f}_{H}$).
The f{f}-harmonic forms have similar properties as in the classical case, e.g., [8]. Let (37) and (15) hold on a closed manifold $(M,G=g+F)$, then for $\omega\in\Lambda^{k}(M)$, by Proposition 5, we have
$({\mathrm{\Delta}}^{f}_{H}\,\omega,\,\omega)_{L^{2}}=(d^{f}\omega,\,d^{f}\omega)_{L^{2}}+(\breve{\delta}^{\,{f}}\omega,\,\breve{\delta}^{\,{f}}\omega)_{L^{2}};$
thus, such $\omega$ is f{f}-harmonic (and similarly for $\breve{\mathrm{\Delta}}^{f}_{H}$-harmonic case) if and only if $d^{f}\omega=0$ and $\breve{\delta}^{\,{f}}\omega=0$. If ${\mathrm{\Delta}}^{f}_{H}\,\omega=0$ and $\omega=d^{f}\,\theta$ hold for $\omega\in\Lambda^{k}(M)$, then $\breve{\delta}^{\,{f}}d^{f}\theta=\breve{\delta}^{\,{f}}\omega=0$; thus,
$(\omega,\,\omega)_{L^{2}}=(d^{f}\theta,\,d^{f}\theta)_{L^{2}}=(\theta,\,\breve{\delta}^{\,{f}}d^{f}\theta)_{L^{2}}=(\theta,\,\breve{\delta}^{\,{f}}\omega)_{L^{2}}=0.$
Moreover, if $\omega\in\Lambda^{k}(M)$ is f{f}-harmonic and $\omega=d^{f}\,\theta$ for some $\theta\in\Lambda^{k-1}(M)$, then $\omega=0$.
@@
Similarly, we obtain the second equality in (47). From this and (32) the equalities (48) follow. ∎
We will modify Stokes theorem on a complete open Riemannian manifold, see also Theorem 1.
@@
We call ${\mathrm{\Delta}}^{f}\psi=\operatorname{div}_{f}(\nabla^{f}\psi)$ the f{f}-Laplacian for a function $\psi\in C^{2}(M)$. By (22), we obtain
${\mathrm{\Delta}}^{f}\psi={{\mathrm{\Delta}}}^{0}\psi+({f}E)(\psi).$
The next theorem extends the well-known classical result on subharmonic functions.
@@
Let a manifold $(M,G=g+F)$ be complete noncompact, and (36) hold for an Einstein’s connection $\nabla$. Assume that $\|{f}\nabla^{f}\psi\|_{g}$ and $\|\psi{f}\nabla^{f}\psi\|_{g}$ belong to $L^{1}(M,g)$ for some $\psi\in C^{2}(M)$ satisfying either ${\mathrm{\Delta}}^{f}\psi\geq 0$ or ${\mathrm{\Delta}}^{f}\psi\leq 0$. Then $\nabla^{f}\psi=0$; moreover, if either $f(TM)=TM$ (hence, $\dim M$ is even), or $f(TM)\neq TM$ and f(TM){f}(TM) is bracket-generating, then $\psi\!=\operatorname{const}$.
@@
Set $X=\nabla\,^{f}\psi$, then $\Delta^{f}\,\psi=\operatorname{div}_{f}X$. By Proposition 6 with $X=\nabla\,^{f}\psi$ and condition $\|{f}\nabla^{f}\psi\|\in L^{1}(M,g)$, we get $\Delta\,^{f}\,\psi\equiv 0$. Using
$\operatorname{div}_{f}(\psi\cdot Y)=\psi\cdot\operatorname{div}_{f}\,Y+g(\nabla^{f}\psi,\ Y)$
with $Y=\nabla\,^{f}\psi$, Proposition 6 with $X=\psi\nabla^{f}\psi$ and condition $\|\psi\,{f}\nabla\,^{f}\psi\|\in L^{1}(M,g)$, we get $(\nabla^{f}\psi,\,\nabla^{f}\psi)_{L^{2}}=0$, hence $\nabla^{f}\psi=0$. If $f(TM)=TM$, then $\psi\!=\operatorname{const}$. If $f(TM)\neq TM$, then $\nabla^{f}\psi=0$ means that $\psi\in C^{2}(M)$ is constant along the integral curves of $f(TM)$; moreover, if the singular distribution $f(TM)$ is bracket-generating then, by Chow’s theorem [2], $\psi={\rm const}$. ∎

### v2

@@ 3 Main results
Here, motivated by the almost Lie algebroid structure, we develop Bochner’s technique for NGT spaces. In Example 3, using the condition (27), we explicitly obtain an Einstein’s connection.
@@ 3.1 The modified covariant derivative and bracket
and the difference tensor $K$, see (8), but generally $\nabla^{f}$ and $\nabla^{0}$ are not linear connections on $M$. Set $\nabla^{f}_{X}\,\psi=({f}X)\psi=d\psi({f}X)$ for the f{f}-gradient of $\psi\in C^{1}(M)$. The f{f}-derivative of an $(p,k)$-tensor $S$, where $p=0,1$, is the following $(p,k+1)$-tensor:
$(\nabla^{f}S)(Y,X_{1},\ldots,X_{k}){=}\nabla^{f}_{Y}(S(X_{1},\ldots,X_{k})){-}\sum\nolimits_{\,i=1}^{k}S(X_{1},\ldots,\nabla^{f}_{Y}X_{i},\ldots,X_{k}).$ (23)
Such $\nabla^{f}$ (as well as $\nabla^{0}$) satisfies Koszul conditions, see (21),
@@
We use the standard notation $\nabla^{f}_{Y}\,S=(\nabla^{f}S)(Y,\ldots)$. A tensor $S$ is called $\nabla^{f}$-parallel if $\nabla^{f}S=0$. Using (23) and $\nabla^{g}\,g=0$, we obtain the following:
@@
Example 3.
If condition (27) is true, then we obtain an almost Lie algebroid structure on $M$ with an anchor $f:TM\to M$ and a skew symmetric bracket $\cdot,\,\cdot\,]_{f}$ on ${\cal E}=TM$ satisfying (18). On the other hand, from (27), we obtain the torsion $T$, hence, an example of the Einstein’s connection $\nabla$.
For an f{f}-connection $\nabla^{f}$ on $(M,G=g+F)$, its dual f{f}-connection $\breve{\nabla}^{f}$ is defined by
@@
If $k=1$, the last term in the above does not appear. If $k>1$, we fix $2\leq p\leq k$ and find
@@ 3.2 The f{f}-divergence and Bochner f{f}-Laplacian
using a local orthonormal frame $\{e_{i}\}$. The divergence of a vector field $X$ can also be represented as follows:
@@
The “musical” isomorphisms $\sharp:T^{*}M\to TM$ and $\flat:TM\to T^{*}M$ are used for rank one tensors, e.g. if $\omega\in\Lambda^{1}(M)$ and $X\in{\mathfrak{X}}_{M}$ then $\omega(X)=g(\omega^{\sharp},X)=g(\omega,X^{\flat})=X^{\flat}(\omega^{\sharp})$. In order to generalize the Stokes Theorem, we formulate the following.
@@
where $\nu$ is the unit inner normal to $\partial M$. On a manifold $(M,G=g+F)$ without boundary, for any $X\in\mathfrak{X}_{M}$ with compact support, (38) reduces to the equality $\int_{M}(\operatorname{div}_{f}X)d\operatorname{vol}_{g}{=}0$.
@@
while, for $k$-forms, we set $g(\omega_{1},\,\omega_{2})=\sum\nolimits_{\,i_{1}<\ldots\,<i_{k}}\omega_{1}(e_{i_{1}},\ldots,\,e_{i_{k}})\,\omega_{2}(e_{i_{1}},\ldots,\,e_{i_{k}})$.
@@
The $L^{2}$-adjoint operator $\nabla^{*{f}}$ (and similarly $\breve{\nabla}^{*{f}}$) for $\nabla^{f}$ maps $(p,k+1)$-tensors, where $p=0,1$, to $(p,k)$-tensors:
$(\nabla^{*{f}}S)(X_{1},\ldots,X_{k})=-\sum\nolimits_{\,i}(\nabla^{f}_{e_{i}}S)(e_{i},X_{1},\ldots,X_{k}).$ (39)
@@
Define a compactly supported 1-form $\omega$ by $\omega(Y)=g(\iota_{\,Y}\,\omega_{2},\,\omega_{1})$, where $Y\in{\mathcal{X}}_{M}$. To simplify calculations, assume that $k=1$. Take an orthonormal frame $(e_{i})$ such that $\nabla_{Y}e_{i}=0$ for $Y\in T_{x}M$. To simplify calculations, assume that $s=t=1$, then at $x\in M$,
$-\nabla^{*{f}}\omega=\sum\nolimits_{j}(\nabla^{f}_{e_{j}}\omega)(e_{j})=\sum\nolimits_{i,j}g({f}e_{j},e_{i})\,(\nabla_{e_{i}}\omega)(e_{j})$
$=\sum\nolimits_{i,j,c}g({f}e_{j},e_{i})\,\nabla_{e_{i}}g(\omega_{2}(e_{j},e_{c}),\omega_{1}(e_{c})\,)$
@@
The operator $\nabla^{*{f}}\,\nabla^{f}$ will be called the Bochner f{f}-Laplacian for an f{f}-connection.
@@
Let an Einstein’s connection $\nabla$ on a closed manifold $(M,G=g+F)$ satisfy (15) and (36). If $g(\nabla^{*f}\,\nabla^{f}\omega,\,\omega)\leq 0$ holds for a form $\omega\in\Lambda^{k}(M)$, then $\nabla^{f}\omega=0$.
@@
hence, $\omega$ is f{f}-parallel: $\nabla^{f}\omega=0$. ∎
@@ 3.3 The Hodge f{f}-Laplacians
Using an Einstein’s connection, define the exterior f{f}-derivative of $\omega\in\Lambda^{k}(M)$ by
@@
The f{f}-derivative $d^{f}:{\mathrm{\Omega}}^{k}(M)\rightarrow{\mathrm{\Omega}}^{k+1}(M)$ is a 1-degree derivation, see (2.2):
@@
+ $+\sum\limits_{\,0\leq i<j\leq k}(-1)^{i+j}\omega\big([X_{i},X_{j}]_{f},X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{j-1},X_{j+1},\ldots,X_{k}\big).$ (45)
@@
+ $+\sum\limits_{\,0\leq i<j\leq k}(-1)^{i+j}\omega\big([X_{i},X_{j}]^{\smile}_{f},X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{j-1},X_{j+1},\ldots,X_{k}\big).$ (46)
@@
+ $+\sum\limits_{\,0\leq i<j\leq k}(-1)^{i+j}\omega\big(\nabla^{f}_{X_{i}}X_{j}-\nabla^{f}_{X_{j}}X_{i},X_{0},\ldots,X_{i-1},X_{i+1},\ldots,X_{j-1},X_{j+1},\ldots,X_{k}\big).$
@@
Define the f{f}-codifferential $\delta^{f}:\Lambda^{k}(M)\to\Lambda^{k-1}(M)$ by $\delta^{\,f}:=\nabla^{*{f}}$. Similarly, we set $\breve{\delta}^{\,f}:=\breve{\nabla}^{*{f}}$. Using the analog of (39), we obtain
@@
On a closed manifold $(M,G=g+F)$ with an Einstein’s connection satisfying (15) and (36), for any $\omega_{1}\in\Lambda^{k}(M)$ and $\omega_{2}\in\Lambda^{k+1}(M)$, we have $(\breve{\delta}^{\,{f}}\omega_{2},\,\omega_{1})_{L^{2}}=(\omega_{2},\,d^{f}\omega_{1})_{L^{2}}$.
@@
$g(d^{f}\omega_{1},\,\omega_{2})=$
$=\sum\nolimits_{\,u=0}^{k}(-1)^{i}\,\nabla^{f}_{\partial_{i_{u}}}\omega_{1}(\partial_{i_{0}},\ldots,\partial_{i_{u-1}},\partial_{i_{u+1}},\ldots,\partial_{i_{k}})g^{i_{0}j_{0}}\ldots g^{i_{k}j_{k}}\omega_{2}(\partial_{j_{0}},\ldots,\partial_{j_{k}})$
@@
A form $\omega$ is called f{f}-harmonic if ${\mathrm{\Delta}}_{H}^{\,{f}}\,\omega=0$ and $\|\omega\|_{L^{2}}<\infty$ (similarly for $\breve{\mathrm{\Delta}}^{f}_{H}$).
The f{f}-harmonic forms have similar properties as in the classical case, e.g., [8]. Let (37) and (15) hold on a closed manifold $(M,G=g+F)$, then for $\omega\in\Lambda^{k}(M)$, by Proposition 5, we have $({\mathrm{\Delta}}^{f}_{H}\,\omega,\,\omega)_{L^{2}}=(d^{f}\omega,\,d^{f}\omega)_{L^{2}}+(\breve{\delta}^{\,{f}}\omega,\,\breve{\delta}^{\,{f}}\omega)_{L^{2}}$; such $\omega$ is f{f}-harmonic (and similarly for $\breve{\mathrm{\Delta}}^{f}_{H}$-harmonic case) if and only if $d^{f}\omega=0$ and $\breve{\delta}^{\,{f}}\omega=0$. If ${\mathrm{\Delta}}^{f}_{H}\,\omega=0$ and $\omega=d^{f}\,\theta$ hold for $\omega\in\Lambda^{k}(M)$, then $\breve{\delta}^{\,{f}}d^{f}\theta=\breve{\delta}^{\,{f}}\omega=0$; thus,
$(\omega,\,\omega)_{L^{2}}=(d^{f}\theta,\,d^{f}\theta)_{L^{2}}=(\theta,\,\breve{\delta}^{\,{f}}d^{f}\theta)_{L^{2}}=(\theta,\,\breve{\delta}^{\,{f}}\omega)_{L^{2}}=0;$
and if $\omega\in\Lambda^{k}(M)$ is f{f}-harmonic and $\omega=d^{f}\,\theta$ for some $\theta\in\Lambda^{k-1}(M)$, then $\omega=0$.
@@
Similarly, we get the second equality in (47). This and (32) yield the equalities (48). ∎
We will modify Stokes theorem on a complete open Riemannian manifold.
@@
We call ${\mathrm{\Delta}}^{f}\psi=\operatorname{div}_{f}(\nabla^{f}\psi)$ the f{f}-Laplacian for a function $\psi\in C^{2}(M)$. By (22), ${\mathrm{\Delta}}^{f}\psi={\mathrm{\Delta}}^{0}\psi+({f}E)(\psi)$.
The next theorem extends the classical result on subharmonic functions.
@@
Let a manifold $(M,G=g+F)$ be complete noncompact, and (36) hold for an Einstein’s connection $\nabla$. If $\|{f}\nabla^{f}\psi\|_{g}$ and $\|\psi{f}\nabla^{f}\psi\|_{g}$ belong to $L^{1}(M,g)$ for some $\psi\in C^{2}(M)$ satisfying either ${\mathrm{\Delta}}^{f}\psi\geq 0$ or ${\mathrm{\Delta}}^{f}\psi\leq 0$, then $\nabla^{f}\psi=0$; moreover, if either $f(TM)=TM$ ($\dim M$ is even), or $f(TM)\neq TM$ and f(TM){f}(TM) is bracket-generating, then $\psi\!=\operatorname{const}$.
@@
Set $X=\nabla\,^{f}\psi$, then $\Delta^{f}\,\psi=\operatorname{div}_{f}X$. By Proposition 6 with $X=\nabla\,^{f}\psi$ and condition $\|{f}\nabla^{f}\psi\|\in L^{1}(M,g)$, we get $\Delta\,^{f}\,\psi\equiv 0$. Using $\operatorname{div}_{f}(\psi\cdot Y)=\psi\cdot\operatorname{div}_{f}\,Y+g(\nabla^{f}\psi,\ Y)$ with $Y=\nabla\,^{f}\psi$, Proposition 6 with $X=\psi\nabla^{f}\psi$ and condition $\|\psi\,{f}\nabla\,^{f}\psi\|\in L^{1}(M,g)$, we get $(\nabla^{f}\psi,\,\nabla^{f}\psi)_{L^{2}}=0$, hence $\nabla^{f}\psi=0$. If $f(TM)=TM$, then $\psi\!=\operatorname{const}$. If $f(TM)\neq TM$, then $\nabla^{f}\psi=0$ means that $\psi$ is constant along the curves in $f(TM)$; moreover, if the singular distribution $f(TM)$ is bracket-generating then, by Chow’s theorem [2], $\psi={\rm const}$. ∎